5. What is the yield to maturity of a consol with a coupon of $85 and a price of $1,000? A) 5.56% B) 8.50% C) 9.00% D) Not enough information has been provided to determine the answer 6. A college graduate has a student loan of $23,000. It is scheduled to be paid back in 10 years with a payment starts next month. If each monthly payment total of 120 monthly payments, and the first is $226.36, what is the interest rate (yield to maturity) on this loan? (Hint: use the fixed payment loan calculator provided in Notes 2 on page 12. The calculator can be used for any fixed payment loan.) A) 0.98% B) 3.40% C) 5.42% D) 6.60% 7. Which of the following represents the equation that would be used to determine the yield to maturity of a three-year fixed payment loan of $1400 which has annual payments of $500 per year? A) $1400 $500/(1i)3 В) і ($1400-5500)/$1400 C) $1400 $500/(1i)$500/(1i$500/(1i)3 D) $1400 $500/(1i) $500/(1i)2 $500(1i)3$1400/(1i)3 8. If, while you are holding a coupon bond, the interest rates on other similar bonds fall, you can be sure that A) the coupon payments on your bond will fall. B) the market price of your bond will rise C) the market price of your bond will fall. D) the par value of your bond will rise
Answer #1

5.

Yeild to maturity of consol = 85 / 1000 = 0.085 = 8.5%

Option B is correct

6.

Loan amount = 23000, tenure = 120 monthly payments, monthly payments = 226.36

let monthly yeild to maturity be i, then

23000 = 226.36 * ( P/A, i%, 120)

( P/A, i%, 120) = 23000 / 226.36 = 101.6081

using trail and error method

when i = 0.25% , value of ( P/A, i%, 120) = 103.5618

when i = 0.5% , value of ( P/A, i%, 120) = 90.07345

Using interpolation

i = 0.25% + (103.5618 - 101.6081) / (103.5618 - 90.07345) * (0.5% - 0.25%)

i = 0.25% + 0.03621%

i = 0.2862%

Yeild to maturity = (1+0.002862)^12 -1 = 1.0348911 - 1 = 0.0348911 = 3.48%

Option B is correct, minor diff due to hand calculations and in between steps being decimal calculations

7

P = F / (1+i)^t

Loan amount = 1400, payment per year for three years = 500

Let yeild to maturity be i

Then,

1400 = 500/(1+i) + 500/(1+i)^2 + 500/(1+i)^3

Option C is correct

8.

if market interest falls, then the market price of bonds increase

So option B is correct